# On the Characteristics of the Conjugate Function Enabling Effective Dual Decomposition Methods

### v1

@@ V-C Problem (30) with $\ell_{\infty}$-norm Constraints
In this section, we exploit specific structural properties of the constraint set $\mathcal{Z}$ in the consensus problem (29) to yield feasible $\mathbf{\bar{y}}\in\mathcal{Y}$ for the feasibility problem (26), where $\mathcal{Y}=\mathcal{Z}^{m}$. In particular, when $\mathcal{Z}$ conforms to $\ell_{\infty}$-norm constraints  [^3] Typically, $\ell_{\infty}$-norm constraints appear in machine learning applications. For example, the $l_{\infty}$ regularization can be considered as a form of constraint on the $l_{\infty}$ norm of parameters., we show that problem (26) associated to (30) is feasible for some particular $\mathbf{\bar{y}}\in\mathcal{Y}$. We first lay out the following remark which will be useful in the sequel for our exposition.
@@
Let Assumption 1 and Assumption 2 hold. Suppose $\mathcal{Z}=\{\mathbf{z}\in{\rm I\!R}^{n}\ |\ \|\mathbf{z}\|_{\infty}\leq a\}$, where $a\in{\rm I\!R}$. Let $\bar{\mathbf{y}}=[\bar{\mathbf{y}}_{1}^{\mbox{\scriptsize T}}\ \ldots\ \bar{\mathbf{y}}_{m}^{\mbox{\scriptsize T}}]^{\mbox{\scriptsize T}}\in\texttt{bnd}~{}\mathcal{Z}^{m}$, where $\forall i\in\mathcal{S}$,
@@
Then problem (26) associated to problem (30) is feasible.
@@
It is straightforward to see that the unique solutions to the feasibility problem (26) associated to problem (30) are given by $\bm{\tilde{\mu}}=(1/\beta)\mathbf{A}\mathbf{\bar{y}}$ and $\bm{\tilde{\eta}}=\mathbf{A}^{\mbox{\scriptsize T}}\bm{\tilde{\mu}}$. Thus, with our choice of $\mathbf{\bar{y}}$, we have
@@
and $\text{and }\bm{\tilde{\eta}}{=}\begin{cases}\displaystyle(1/\beta)[2a\mathbf{1}_{n}^{\mbox{\scriptsize T}}\ -2a\mathbf{1}_{n}^{\mbox{\scriptsize T}}]^{\mbox{\scriptsize T}}\ ;&\text{$m=2$}\\ \displaystyle(1/\beta)[2a\mathbf{1}_{n}^{\mbox{\scriptsize T}}\ -4a\mathbf{1}_{n}^{\mbox{\scriptsize T}}\ 4a\mathbf{1}_{n}^{\mbox{\scriptsize T}}\ -4a\mathbf{1}_{n}^{\mbox{\scriptsize T}}\ \ldots\ 2a\mathbf{1}_{n}^{\mbox{\scriptsize T}}]^{\mbox{\scriptsize T}};&\\ \ \hskip 113.81102pt\text{$m$ odd, $m\geq 3$}\\ \displaystyle(1/\beta)[2a\mathbf{1}_{n}^{\mbox{\scriptsize T}}\ -4a\mathbf{1}_{n}^{\mbox{\scriptsize T}}\ 4a\mathbf{1}_{n}^{\mbox{\scriptsize T}}\ -4a\mathbf{1}_{n}^{\mbox{\scriptsize T}}\ \ldots\ -2a\mathbf{1}_{n}^{\mbox{\scriptsize T}}]^{\mbox{\scriptsize T}};&\\ \ \hskip 113.81102pt\text{$m$ even, $m\geq 4$}.\end{cases}$ (38)
Clearly, $\bm{\tilde{\eta}}\in N_{\mathcal{Y}}(\mathbf{\bar{y}})$, where $\mathcal{Y}=\mathcal{Z}^{m}$, cf.  Remark 5. Thus $\bm{\tilde{\eta}}\in\partial\delta_{\mathcal{Y}}(\mathbf{\bar{y}})$, cf.  Remark 2. Hence problem (26) is feasible. ∎
Lemma 1 indicates that problem (26) associated to problem (30) is feasible if we select $\mathbf{\bar{y}}$ according to (36), and the unique solution $(\tilde{\bm{\eta}},\tilde{\bm{\mu}})$ can be derived. Thus, as we have already pointed out in Section V-B, we can choose $\bm{\lambda}^{(0)}{=}\bm{\tilde{\mu}}$ for some $\beta{\in}{\rm I\!R}$ [cf.  (37)] as an initialization point of Algorithm 3. In the sequel, we numerically evaluate the performance of Algorithm 3 with this choice of initialization point.
@@ VI Numerical Results
In this section, we numerically evaluate the benefits of FGOR characteristics for solving the global consensus problem (29), cf.  § V. In particular, we compare the performance of Algorithm 3 against Algorithm 2. The comparisons are demonstrated with the following stepsize rules: 1) traditional stepsizes: constant stepsize $\gamma_{k}=\gamma_{0}$, polynomially decay stepsizes $\gamma_{k}=\gamma_{0}/k,\gamma_{0}/\sqrt{k}$, and geometrically decay stepsize $\gamma_{k}=\gamma_{0}q^{k}$, where $\gamma_{0}>0$ is an appropriately chosen constant, 2) step decay stepsize [23], 3) stepsize rule [35]. Moreover, the results are illustrated under two cases, Case 1: the local functions $f_{i}$s in the distributed consensus formulation (30) are quadratic, and Case 2: problem (30) is the regularized least squares regression with a real data set. For each case, we conduct experiments for $\mathbf{y}^{\star}\in\texttt{int}~{}\mathcal{Y}$ and $\mathbf{y}^{\star}\in\texttt{bnd}~{}\mathcal{Y}$, where $\mathcal{Y}=\mathcal{Z}^{m}$ and $\mathcal{Z}$ is an $\ell_{\infty}$-norm constraint, cf.  Lemma 1. For Algorithm 3, we choose $\bm{\lambda}^{(0)}=\bm{\tilde{\mu}}$ for some $\beta\in{\rm I\!R}$, cf.  (37). Furthermore, the constant stepsize $\gamma$ of Algorithm 3 [cf.  step 7] is chosen to be sufficiently small to avoid undesired stepping over the region $\mathcal{V}$, cf.  (3). Note that the initialization point $\bm{\tilde{\mu}}$ of Algorithm 3 is deterministic and is computed in advance, cf.  problem (26), (37). However, since the initialization for Algorithm 2 is arbitrary, for a fair comparison, an average performance of the algorithm is considered. More specifically the initialization point $\bm{\lambda}^{(0)}$ of Algorithm 2 is chosen uniformly over the sphere centered at $\bm{\lambda}^{\star}$ with radius $\|\bm{\tilde{\mu}}-\bm{\lambda}^{\star}\|_{2}$.
Figure 5: Case 1: $\mathbf{y}^{\star}\in\texttt{int}~{}\mathcal{Y}$: Comparison of results with traditional stepsizes, step decay stepsize (SD), and Ref. [35]. LABEL:sub@subfig:linear-regression-1 Traditional stepsizes: $m=4$ and $n=3$. LABEL:sub@subfig:linear-regression-2 Traditional stepsizes: $m=4$ and $n=10$. LABEL:sub@subfig:linear-regression-3 SD and Ref. [35]: $m=4$ and $n=3$. LABEL:sub@subfig:linear-regression-4 SD and Ref. [35]: $m=4$ and $n=10$.
VI-A Case 1: Experiments with Quadratic Local Objectives
@@
We consider problem (30) with quadratic $f_{i}$s, i.e., $f_{i}(\mathbf{y}_{i})=\mathbf{y}_{i}^{\mbox{\scriptsize T}}\mathbf{A}_{i}\mathbf{y}_{i}+\mathbf{q}_{i}^{\mbox{\scriptsize T}}\mathbf{y}_{i},\,i\in\mathcal{S}$, where $\mathbf{A}_{i}\in\mathbb{S}^{n}_{++}$ and $\mathbf{q}_{i}\in{\rm I\!R}^{n}$ are arbitrarily chosen.
@@
Figure 5 and Figure 6 show the comparison of results when $\mathbf{y}^{\star}\in\texttt{int}~{}\mathcal{Y}$ and $\mathbf{y}^{\star}\in\texttt{bnd}~{}\mathcal{Y}$, respectively. In particular, Figure 5 shows the convergence of dual function values using $m=4$ agents when the lengths of $\mathbf{y}_{i}$s are $n=3$ and $n=10$. Figure 6 captures the corresponding convergences with $m=3$ and $m=5$ using scalar-valued $\mathbf{y}_{i}$s. For the constant stepsize rule, we choose $\gamma_{k}=1/L_{g}$, where $L_{g}$ represents the gradient Lipschitz constant of $g$. Both figures show that Algorithm 3 achieves significantly faster convergences than Algorithm 2. To further clarify, the comparison of results in Figure 5LABEL:sub@subfig:linear-regression-8 is summarized in Table I. For each method, the number of iterations and the total number of bits required to achieve an accuracy of $10^{-5}$ are given by $k^{\star}$ and $b^{\star}$, respectively. Furthermore, for our method $b^{\star}=k_{0}(m+1)+n(k^{\star}-k_{0})(2m-1)b$ and for the other methods $b^{\star}=nk^{\star}(2m-1)b$, where $k_{0}$ is the number of iterations that $\bm{\lambda}^{(k)}$ lies in the FGOR region and $b$ is the bits used to represent each component of communicated vectors $\mathbf{y}_{i}^{(k)}$ and $\bm{\lambda}_{i}^{(k)}$, cf.  § V-B2.
@@
Figure 6: Case 1: $\mathbf{y}^{\star}\in\texttt{bnd}~{}\mathcal{Y}$: Comparison of results with traditional stepsizes, step decay stepsize (SD), and Ref. [35]. LABEL:sub@subfig:linear-regression-1 Traditional stepsizes: $m=3$ and $n=1$. LABEL:sub@subfig:linear-regression-2 Traditional stepsizes: $m=5$ and $n=1$. LABEL:sub@subfig:linear-regression-3 SD and Ref. [35]: $m=3$ and $n=1$. LABEL:sub@subfig:linear-regression-4 SD and Ref. [35]: $m=5$ and $n=1$.
VI-B Case 2: Regularized Least Squares Regression
@@
We consider problem (30) to be the least squares regression with $\ell_{\infty}$-norm regularization. Then the corresponding regularized problem has the form
$\begin{array}[]{ll}\mbox{minimize}&f_{0}(\mathbf{y})=\sum_{i{=}1}^{m}f_{i}(\mathbf{y}_{i})+d\|\mathbf{y}\|_{\infty}^{2}\\ \mbox{subject to}&\mathbf{y}\in\mathcal{Y}\\ &\mathbf{Ay}=\mathbf{0},\end{array}$ (39)
@@
where $\mathbf{y}=[\mathbf{y}_{1}^{\mbox{\scriptsize T}}\ \ldots\ \mathbf{y}_{m}^{\mbox{\scriptsize T}}]^{\mbox{\scriptsize T}}$, $\mathcal{Y}=\mathcal{Z}^{m}$, $\mathbf{A}$ is given in (33), and $d>0$ is the regularization parameter. As we have already pointed out in Assumption 1 [cf.  § II], the problem of convex regularized minimization is equivalent to its corresponding regularization-constrained formulation [47, § 1.2]. Thus, problem (39) can be equivalently reformulated as
$\begin{array}[]{ll}\mbox{minimize}&f_{0}(\mathbf{y})=\sum_{i{=}1}^{m}f_{i}(\mathbf{y}_{i})\\ \mbox{subject to}&\|\mathbf{y}\|_{\infty}^{2}\leq\bar{d}\\ &\mathbf{Ay}=\mathbf{0},\end{array}$ (40)
where $\bar{d}>0$ is a hyperparameter. We consider each $f_{i}$, $i\in\mathcal{S}$, in (40) with form
$f_{i}(\mathbf{y}_{i})=\frac{1}{N}(\mathbf{y}_{i}^{\mbox{\scriptsize T}}\mathbf{x}_{i}-t_{i})^{2},$ (41)
where $\mathbf{x}_{i}\in{\rm I\!R}^{n}$ is the feature vector, $t_{i}$ is the observation, and $N$ is the number of training examples of agent $i$. We have performed the experiments with the real estate valuation dataset [54] which contains 276 training data samples and 138 testing data samples. The length of the feature vector $\mathbf{x}_{i}$ is $n=6$. Since the features have different units, the data set is standardized before we run the experiments.
@@
Figure 7 and Figure 8 show the comparison of results when $\mathbf{y}^{\star}\in\texttt{int}~{}\mathcal{Y}$ and $\mathbf{y}^{\star}\in\texttt{bnd}~{}\mathcal{Y}$, respectively. Both figures show the convergence of dual function values with $m=3$ and $m=6$, that is, each agent having $N_{1}=46$ and $N_{2}=96$ data samples, respectively. Similarly to Case 1, both figures show that Algorithm 3 converges towards the optimal solution much faster than Algorithm 2. A summary of results similar to Figure 5LABEL:sub@subfig:linear-regression-8 is also provided in Table I for Figure 8LABEL:sub@subfig:linear-regression-4.
@@
Figure 7: Case 2: $\mathbf{y}^{\star}\in\texttt{int}~{}\mathcal{Y}$: Comparison of results with traditional stepsizes, step decay stepsize (SD), and Ref. [35]. LABEL:sub@subfig:linear-regression-5 Traditional stepsizes: $m=3$. LABEL:sub@subfig:linear-regression-6 Traditional stepsizes: $m=6$. LABEL:sub@subfig:linear-regression-7 SD and Ref. [35]: $m=3$. LABEL:sub@subfig:linear-regression-8 SD and Ref. [35]: $m=6$.

### v3

@@ V-C Problem (33) with $\ell_{\infty}$-norm Constraints
In this section, we exploit specific structural properties of the constraint set $\mathcal{Z}$ in problem (33) to provide a point $\mathbf{\bar{y}}\in\texttt{bnd}~\mathcal{Y}$, such that problem (28) associated to (33), is feasible, where $\mathcal{Y}=\mathcal{Z}^{m}$. In particular, when $\mathcal{Z}$ conforms to $\ell_{\infty}$-norm constraints  [^4] Typically, $\ell_{\infty}$-norm constraints appear in machine learning applications. For example, the $l_{\infty}$ regularization can be considered as a form of constraint on the $l_{\infty}$ norm of parameters., we show that problem (28) associated to (33) is feasible for a specific $\mathbf{\bar{y}}\in\mathcal{Y}$. We first present the following remark, which will be useful in the sequel for our exposition.
@@
Let Assumption 1 and Assumption 2 hold. Suppose $\mathcal{Z}=\{\mathbf{z}\in{\rm I\!R}^{n}\ |\ \|\mathbf{z}\|_{\infty}\leq a\}$, where $a\in{\rm I\!R}$. Let $\bar{\mathbf{y}}=[\bar{\mathbf{y}}_{1}^{\mbox{\scriptsize T}}\ \ldots\ \bar{\mathbf{y}}_{m}^{\mbox{\scriptsize T}}]^{\mbox{\scriptsize T}}$, where $\forall i\in\mathcal{S}$,
@@
Then problem (28) associated to problem (33), is feasible.
@@
Clearly, $\mathbf{\bar{y}}\in\texttt{bnd}~\mathcal{Y}$, where $\mathcal{Y}=\mathcal{Z}^{m}$. By solving the equality constraints in (28), we uniquely obtain the solutions $\bm{\tilde{\mu}}=(1/\beta)\mathbf{A}\mathbf{\bar{y}}$ and $\bm{\tilde{\eta}}=\mathbf{A}^{\mbox{\scriptsize T}}\bm{\tilde{\mu}}$. Thus, we have
@@
[…]
Clearly, $\bm{\tilde{\eta}}\in N_{\mathcal{Y}}(\mathbf{\bar{y}})$, cf. Remark 6. Thus, problem (28) associated to problem (33), is feasible. ∎
Note that the feasibility of problem (28) associated to the consensus problem (33), confirms the satisfaction of Assumption 3 for (33), cf.  Remark 3. In the sequel, we numerically evaluate the performance of Algorithm 3.
@@ VI Numerical Results
In this section, we numerically evaluate the benefits of FGOR for solving the consensus problem (33), cf.  § V. In particular, we compare the performance of Algorithm 3 with Algorithm 2 and splitting methods [condat_2013_splitting_methods, Boyd-Parikh-Chu-Peleato-Eckstein-2010, He_2021, Bai_2024]. The experiments are conducted for $\mathcal{Y}=\mathcal{Z}^{m}$ and $\mathcal{Z}$ is defined by an $\ell_{\infty}$-norm constraint, cf.  Lemma 2. For Algorithm 3, we choose $\bm{\lambda}^{(0)}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}c}}\bm{\tilde{\mu}}$, where $c\in{\rm I\!R}$, for some $\beta\in{\rm I\!R}$, cf.  (40). Furthermore, the constant stepsize $\gamma$ in Algorithm 3 [cf.  step 7] is chosen to be sufficiently small to avoid undesired stepping over the region $\mathcal{V}$, cf.  (3). Note that the initialization point ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}c}}\bm{\tilde{\mu}}$ of Algorithm 3 is deterministic and is computed in advance, cf.  problem (28), (40). However, since the initializations for Algorithm 2 and [condat_2013_splitting_methods, Boyd-Parikh-Chu-Peleato-Eckstein-2010] are arbitrary, for a fair comparison, an average performance of the algorithms is considered. Specifically, the initialization point $\bm{\lambda}^{(0)}$ of Algorithm 2 and [condat_2013_splitting_methods] is chosen uniformly over the sphere centered at $\bm{\lambda}^{\star}$ with radius $\|\bm{\tilde{\mu}}-\bm{\lambda}^{\star}\|_{2}$. For [Boyd-Parikh-Chu-Peleato-Eckstein-2010], $\bm{\lambda}^{(0)}$ is chosen uniformly over the sphere centered at $\bm{\lambda}_{\texttt{admm}}^{\star}$ with radius $\|\bm{\tilde{\mu}}-\bm{\lambda}^{\star}\|_{2}$, where $\bm{\lambda}_{\texttt{admm}}^{\star}$ is the dual optimal point of the corresponding dual problem of [Boyd-Parikh-Chu-Peleato-Eckstein-2010].
VI-A Comparisons with Algorithm 2
The comparisons with Algorithm 2 are demonstrated with the following stepsize rules: 1) traditional stepsizes (TS): constant stepsize $\gamma_{k}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}1/L_{g}}}$, where $L_{g}$ represents the gradient Lipschitz constant of $g$, polynomially decay stepsizes $\gamma_{k}=\gamma_{0}/k,\gamma_{0}/\sqrt{k}$, and geometrically decay stepsize $\gamma_{k}=\gamma_{0}q^{k}$, where $\gamma_{0}>0$ is an appropriately chosen constant, 2) step decay stepsize (SD)[wang_2021], 3) stepsize rule [Yura_2020]. Moreover, the corresponding results are illustrated under two cases, Case 1: quadratic program formulation, and Case 2: regularized least squares regression with two real datasets.
VI-A1 Case 1: Quadratic Program Formulation
@@
We consider problem (33) with quadratic $f_{i}$s, i.e., $f_{i}(\mathbf{y}_{i})=(1/m)(\mathbf{y}_{i}^{\mbox{\scriptsize T}}\mathbf{A}_{i}\mathbf{y}_{i}+\mathbf{q}_{i}^{\mbox{\scriptsize T}}\mathbf{y}_{i}),\,i\in\mathcal{S}$, where $\mathbf{A}_{i}\in\mathbb{S}^{n}_{++}$ and $\mathbf{q}_{i}\in{\rm I\!R}^{n}$ are arbitrarily chosen.
@@
Figure 5 and Figure 6 show the comparison of results when $\mathbf{y}^{\star}\in\texttt{int}~\mathcal{Y}$ and $\mathbf{y}^{\star}\in\texttt{bnd}~\mathcal{Y}$, respectively, for different values of $m$ and $n$. Both figures show that Algorithm 3 achieves significantly faster convergences than Algorithm 2. To further clarify, the comparison of results in Figure 5LABEL:sub@subfig:Case1-11 is summarized in Table I. For each method, the number of iterations and the total number of bits required to achieve an accuracy of $10^{-5}$ are given by $k^{\star}$ and $b^{\star}$, respectively. Furthermore, for our method $b^{\star}=k_{0}(m+1)+n(k^{\star}-k_{0})(2m-1)b$ and for the other methods $b^{\star}=nk^{\star}(2m-1)b$, where $k_{0}$ is the number of iterations that $\bm{\lambda}^{(k)}$ lies in the RFGOR region and $b$ is the bits used to represent each component of communicated vectors $\mathbf{y}_{i}^{(k)}$ and $\bm{\lambda}_{i}^{(k)}$, cf.  § V-B2.
@@
Figure 5: Comparisons with Algorithm 2: Case 1: $\mathbf{y}^{\star}\in\texttt{int}~\mathcal{Y}$. LABEL:sub@subfig:Case1-9 TS: $m=4$ and $n=10$. LABEL:sub@subfig:Case1-10 TS: $m=100$ and $n=10$. LABEL:sub@subfig:Case1-11 SD and Ref. [Yura_2020]: $m=4$ and $n=10$. LABEL:sub@subfig:Case1-12 SD and Ref. [Yura_2020]: $m=100$ and $n=10$.
VI-A2 Case 2: Regularized Least Squares Regression
@@
We consider the following least squares regression problem with $\ell_{\infty}$-norm regularization:
$\begin{array}[]{ll}\mbox{minimize}&f_{0}(\mathbf{y})=(1/m)\sum_{i{=}1}^{m}f_{i}(\mathbf{y}_{i})+d\|\mathbf{y}\|_{\infty}^{2}\\ \mbox{subject to}&\mathbf{Ay}=\mathbf{0},\end{array}$ (41)
@@
where $f_{i}(\mathbf{y}_{i})=(1/N)\sum_{j=1}^{N}(\mathbf{y}_{i}^{\mbox{\scriptsize T}}\mathbf{x}_{j}-t_{j})^{2}$, $i\in\mathcal{S}$, $\mathbf{x}_{j}\in{\rm I\!R}^{n}$ is the feature vector with $1$ in the first entry for the bias term, $t_{j}$ is the observation, and $N$ is the number of training examples of agent $i$. Moreover, $\mathbf{y}=[\mathbf{y}_{1}^{\mbox{\scriptsize T}}\ \ldots\ \mathbf{y}_{m}^{\mbox{\scriptsize T}}]^{\mbox{\scriptsize T}}$, $\mathbf{A}$ is given in (36), and $d>0$ is the regularization parameter. As we have already pointed out in § II in connection with Assumption 1, the problem of convex regularized minimization is equivalent to its corresponding regularization-constrained formulation, cf.  Appendix B. Thus, problem (41) is equivalently reformulated as:
$\begin{array}[]{ll}\mbox{minimize}&f_{0}(\mathbf{y})=(1/m)\sum_{i{=}1}^{m}f_{i}(\mathbf{y}_{i})\\ \mbox{subject to}&\|\mathbf{y}\|_{\infty}^{2}\leq\bar{d}\\ &\mathbf{Ay}=\mathbf{0},\end{array}$ (42)
where $\bar{d}>0$ is a hyperparameter, and experiments are conducted using the real estate valuation dataset (REV) [Yeh-misc_real_estate_valuation] and year prediction million song dataset (MSD) [year_prediction_msd_203]. The first dataset contains $276$ training data samples and $138$ testing data samples, where the length of the feature vector $\mathbf{x}_{j}$ is $n=7$. The latter dataset contains $463715$ training data samples and $51630$ testing data samples, where the length of the feature vector $\mathbf{x}_{j}$ is $n=91$. Both datasets were standardized before running the experiments.
@@
Figure 7 illustrates the convergence of dual function values for two scenarios: $m=6$ with each agent holding $N=46$ samples for the REV dataset [cf.  Figure 7LABEL:sub@subfig:linear-regression-5 and Figure 7LABEL:sub@subfig:linear-regression-7], and $m=100$ with each agent holding $N=4637$ samples for the MSD dataset [cf.  Figure 7LABEL:sub@subfig:linear-regression-6 and Figure 7LABEL:sub@subfig:linear-regression-8]. As in Case 1, the results indicate that Algorithm 3 converges to the optimal solution substantially faster than Algorithm 2.
@@
Figure 6: Comparisons with Algorithm 2: Case 1: $\mathbf{y}^{\star}\in\texttt{bnd}~\mathcal{Y}$. LABEL:sub@subfig:Case1-5 TS: $m=5$ and $n=1$. LABEL:sub@subfig:Case1-6 TS: $m=100$ and $n=100$. LABEL:sub@subfig:Case1-7 SD and Ref. [Yura_2020]: $m=5$ and $n=1$. LABEL:sub@subfig:Case1-8 SD and Ref. [Yura_2020]: $m=100$ and $n=100$.
VI-B Comparisons with Splitting Methods [condat_2013_splitting_methods, Boyd-Parikh-Chu-Peleato-Eckstein-2010, He_2021, Bai_2024]
Let us now compare our proposed Algorithm 3 with the state-of-the-art splitting methods [condat_2013_splitting_methods, Algorithm 3.1], [Boyd-Parikh-Chu-Peleato-Eckstein-2010, § 3.1], [He_2021, § 4.2], and [Bai_2024, § 3]. The comparisons are illustrated when the local functions $f_{i}$s in (33) are quadratic, i.e., of the same form considered in VI-A1. For the comparisons, Algorithm 3 uses the following stepsizes as the switching stepsize rules: 1) constant stepsize $\gamma_{k}=1/L_{g}$, where $L_{g}$ is the gradient Lipschitz constant of $g$, 2) polynomially decay stepsize $\gamma_{k}=\gamma_{0}/k$, where $\gamma_{0}>0$ is an appropriately chosen constant, 3) stepsize rule [Yura_2020]. We note that the dual function formulations considered in [Boyd-Parikh-Chu-Peleato-Eckstein-2010] and in our paper for the primal problem (33) are different. Therefore, instead of comparing the convergence of dual function values, we consider the convergence of primal function values $f_{0}(\mathbf{y}_{\texttt{F}}^{(k)})$, where $\mathbf{y}_{\texttt{F}}^{(k)}=[{\mathbf{y}_{\texttt{f}}}^{(k)\mbox{\scriptsize T}}\,\ldots\,{\mathbf{y}_{\texttt{f}}}^{(k)\mbox{\scriptsize T}}]^{\mbox{\scriptsize T}}$ and $\mathbf{y}_{\texttt{f}}$ is the average of $\mathbf{y}_{i}^{(k)}$s, $i=1,\ldots,m$.
Figure 8LABEL:sub@subfig:Case1-splitting-1 and Figure 8LABEL:sub@subfig:Case1-splitting-2 present a comparison of the results obtained using [condat_2013_splitting_methods, Algorithm 3.1] and [Boyd-Parikh-Chu-Peleato-Eckstein-2010, § 3.1] for cases where $\mathbf{y}^{\star}\in\texttt{int}~\mathcal{Y}$ with $m=4$, $n=3$ and $\mathbf{y}^{\star}\in\texttt{bnd}~\mathcal{Y}$ with $m=5$, $n=1$, respectively. Both figures show that Algorithm 3 outperforms [condat_2013_splitting_methods, Algorithm 3.1] and [Boyd-Parikh-Chu-Peleato-Eckstein-2010, § 3.1] on average. To further highlight the improvements in convergence and communication overhead, a summary of the results for Figure 8LABEL:sub@subfig:Case1-splitting-2 is provided in Table I. Figure 8LABEL:sub@subfig:Case1-splitting-3 and Figure 8LABEL:sub@subfig:Case1-splitting-4 present comparisons with benchmark methods, namely the balanced augmented Lagrangian method (ALM) [He_2021, § 4.2] and the double-penalty ALM [Bai_2024, § 3]. Both figures indicate that the proposed Algorithm 3 outperforms the method in [Bai_2024, § 3]. Moreover, Algorithm 3 demonstrates faster convergence during the initial iterations, however, as the iterations progress, the method in [He_2021, § 4.2] attains better overall convergence performance.
Figure 7: Comparisons with Algorithm 2: Case 2. LABEL:sub@subfig:linear-regression-5 TS: REV dataset with $m=6$. LABEL:sub@subfig:linear-regression-6 TS: MSD with $m=100$. LABEL:sub@subfig:linear-regression-7 SD and Ref. [Yura_2020]: REV dataset with $m=6$. LABEL:sub@subfig:linear-regression-8 SD and Ref. [Yura_2020]: MSD with $m=100$.